A total of 95 theme park tickets were sold for $960. Each adult ticket cost $12 and each child's ticket cost $9.Find the number of adult tickets and the number of children's tickets sold.
Answer by Cromlix(4381)

x is an adults ticket
y is a child's ticket
x + y = 95 .....1
12x + 9y = 960 ....2
Multiply (1) by 9
9x + 9y = 855 ....3
12x + 9y = 960 ....2
Subtract 3 from 2
3x = 105
x = 35
Substitute x = 35 into (1)
x + y = 95
35 + y = 95
y = 95 - 35
y = 60
35*$12 + 60*$9 = $960
Hope this helps
:-)